# IDEAL: Influence-Driven Selective Annotations Empower In-Context Learners in Large Language Models

### v2

@@ 4.3.2 Comparisons with alternative methods
We also compare our method with other alternative methods that can select the coreset from large-scale unlabeled data. We perform the evaluations on MRPC, MNLI and HellaSwag. We include the following alternative methods (1) $K$-Means (Lloyd, 1982), which groups all examples into $m$ clusters, and selects the centroid example from each cluster. (2) Maximizing facility location (MFL) (Lin & Bilmes, 2009), which aims at optimizing the representativeness of the selected subset. (3) Fast Vote-$k$ (Su et al., 2023), which is an efficient alternative to Vote-$k$ which directly picks $m$ examples with the largest Vote-$k$ scores.
@@
Method Datasets Selection Retrieval MRPC MNLI HellaSwag Vote-$k$ Similar 64.6 38.9 67.9 IDEAL Similar 66.4 41.0 68.6 Vote-$k$ Random 60.7 37.8 64.6 IDEAL Random 62.5 39.0 66.8
Method Models Test Domain IMDb BoolQ Cst. Vote-$k$ GPT-Neo 71.1 56.4 IDEAL GPT-Neo 72.2 58.0 Vote-$k$ GPT-J 76.4 56.1 IDEAL GPT-J 76.8 56.4
@@ 4.3.3 Evaluation with different retrieval methods
In previous experiments, we used a similarity-based prompt retrieval method by default. In this section, we conduct experiments to quantify the effect of different prompt retrieval methods under the annotation 100. We present the results in Table 4. We observe that both Vote-$k$ and IDEAL suffer from a significant performance drop when the prompt retrieval method is changed from similarity-based to random selection. Notably, IDEAL also achieves better performance than Vote-$k$ when combined with random retrieval in all datasets. It suggests that IDEAL can cultivate a more stable training subset (Chang & Jia, 2023) for in-context learning tasks. Note that we also show that our IDEAL is more stable and robust against the order of in-context examples in Appendix C.5.
@@ 4.3.4 Evaluation on other language models
Here we evaluate IDEAL on other language models, including GPT-Neo 2.7B (Black et al., 2021), and the advanced chat model GPT-3.5-Turbo where we use the same instruction as other language models for each dataset. While GPT-3.5-Turbo has mainly been optimized for chat, it also performs well on traditional completion tasks (Kheiri & Karimi, 2023). To conduct experiments, we select three classification tasks (MRPC, MNLI, and RTE), considering they are easier for prompting GPT-3.5-Turbo to return responses without pleasantries or explanatory content.
The evaluation results are presented in Figure 5. Our evaluations reveal that IDEAL consistently outperforms the baselines across all models tested. This demonstrates the versatility of our method in the context of learning tasks using models of varying sizes. Notably, we observe that the largest model, i.e., GPT-3.5-Turbo, performs worse than GPT-Neo and GPT-J. This situation arises because GPT-3.5-Turbo is primarily optimized for chat tasks and faces challenges in following human instructions for classification. This scenario also has been identified in Ye et al. (2023).
@@ 4.3.5 Evaluation on out-of-distribution tasks
We further evaluate our method on out-of-distribution tasks (Zhou et al., 2022; Wang et al., 2022b; Zhang et al., 2023b; Huang et al., 2023b; c), where there is a distribution shift between the selective annotation data and test data. Following (Chang & Jia, 2023), we compare IDEAL and Vote-$k$ using SST-2 (Socher et al., 2013), BoolQ (Clark et al., 2019) as source tasks, and IMDb (Maas et al., 2011), BoolQ Contrast Set (Gardner et al., 2020) as target tasks, respectively. In all evaluations, we set the annotation budget as 18 and use the similarity-based retrieve to perform the evaluations on the test set in target domains. We use GPT-J 6B and GPT-Neo 2.7B here and show the results in Table 4. We can observe that IDEAL still outperforms baselines on all datasets with two models, implying that IDEAL could select the subset which could depict the invariant properties of this kind of tasks and generalize to out-of-distribution scenarios.
@@ 4.4 Case study: automatic annotation
In previous experiments, we used a small set of manually annotated examples as candidate prompts to make predictions. In contrast, here we are interested in a case study that utilizes the subset selected by IDEAL to annotate all available unlabeled data automatically, leading to a larger set of candidate prompts. Specifically, we first choose an initial subset from the pool of unlabeled data using IDEAL and manually label this selected subset. Afterward, we simulate the information diffusion process from the initial subset to all other data, where we employ those activated data as prompts to predict upcoming activated data at each step and label them accordingly with prediction results. This process ultimately makes a fully labeled training dataset. Finally, all examples (including manual labeling and automatic labeling ) are utilized as potential prompts in conjunction with the prompt retrieve technique for final testing. We name this paradigm as Auto-IDEAL and compare it with Vote-$k$ and origin IDEAL on all classification datasets. We choose 300 training data for each dataset to perform experiments. The manual annotation budget is set to 150, i.e., half of the labels of the candidate prompts in Auto-IDEAL are annotated automatically. Experimental results are provided in Table 5. As can be observed, Auto-IDEAL even achieves better performance than IDEAL in 4 of 5 cases. Notably, although the performance is worse on MNLI, it is still competitive (better than Vote-$k$). It suggests that expanding the candidate prompts through automatic annotation following the diffusion process can further boost the performance of IDEAL. It benefits from the fact that information only diffuses between similar examples. Therefore, unlabeled examples will be automatically annotated using the most similar annotated examples as prompts leading to a promising annotation success rate.
@@ A.1 In-context learning
In-context learning (ICL) has become a new paradigm for natural language processing (NLP), where large language models make predictions only based on contexts augmented with a few examples (Dong et al., 2023; Xie et al., 2022; Shin et al., 2022; Zhang et al., 2023a; Bai et al., 2023). A series of works attempts to revise, enhance, and understand ICL, which include but are not limited to prompt tuning (Kim et al., 2022; Wang et al., 2022a; Mishra et al., 2022), analyzing intrinsic mechanism (Bansal et al., 2022; Chan et al., 2022; Li et al., 2023a; Garg et al., 2022), evaluations (Srivastava et al., 2023; Wang et al., 2022c), applications in multiple domains (Chen et al., 2022; Lee et al., 2022; Cho et al., 2023), and etc.
Despite in-context learning has shown impressive performance in various domains, its efficacy is sensitive to the selection of in-context learning examples (Zhao et al., 2021b; Lu et al., 2021). Considering this, multiple methods have been proposed to select the optimal in-context learning examples to achieve optimal performance. These methods retrieve the most relevant examples subset with/without specific order for each query (Wu et al., 2022; Liu et al., 2021; Gao et al., 2020) Alternatively, some methods aim to find a set of examples once for all queries on the same task (Li & Qiu, 2023; Diao et al., 2023). Specifically, (Wu et al., 2022) formally defines the problem of self-adaptive In-context learning, which aims to search for the best In-context learning examples and corresponding order for each input query. While, Li & Qiu (2023); Diao et al. (2023) focus on finding one fixed set of examples for each task. However, these methods rely on the assumption that large-scale annotated training examples are always available.
Different from them, this paper studies selective annotations for ICL, which can effectively reduce the annotation cost in ICL. Furthermore, compared with recent work (Su et al., 2023), as discussed in the main paper, this work is superior in many aspects, such as the end-to-end manner, mitigation of the trade-off between diversity and representativeness, theoretical guarantees, and better empirical performance.
@@ A.2 Coreset selection
Coreset selection focuses on selecting a small but highly informative subset from a large dataset for follow-up tasks, which can significantly reduce the data storage cost and training consumption (Huang et al., 2018; 2023a; Feldman & Zhang, 2020; Sorscher et al., 2022; Xia et al., 2023b; Li et al., 2023b; Xia et al., 2022). Most of the works on coreset selection target the scenes of supervised learning and classification (Sener & Savarese, 2018; Toneva et al., 2019; He et al., 2023). Only a few works extend coreset selection into unsupervised cases (Sorscher et al., 2022; Su et al., 2023). This paper studies unsupervised data selection for annotations in ICL, which reduces the annotation expenses of prompts and helps large language models become better few-shot learners. Also, it enjoys theoretical support. Therefore, this work is different from previous efforts and contributes to the research community.
@@ A.3 Data distillation
Data distillation (Wang et al., 2018; Zhao et al., 2021a; Shin et al., 2023; Cui et al., 2023; Du et al., 2023; Loo et al., 2023) is an alternative approach for dataset compression and curation, which is inspired by knowledge distillation. Different from coreset selection, this series of works target synthesizing a small but informative dataset as an alternative to the original dataset. However, data distillation is criticized for only synthesizing a small number of data points due to computational source limitations (Xia et al., 2023a; Yang et al., 2023). The performances of data distillation and data selection are therefore not compared directly. Besides, it is under-explored about how to perform data distillation in an unsupervised manner on natural language processing tasks. Based on this analysis, the data distillation strategy is not involved in empirical evaluations.
@@ B.1.1 Lemma 1
The proof is inspired by (Rolnick & Weed, 2014). We first let
@@ C.1 Selected examples
Dataset
Input
MRPC
a. Input: The two Democrats on the five-member FCC held a press conference to sway opinion against […]
    Output: not equivalent
a. Input: The report shows that drugs sold in Canadian pharmacies are manufactured in facilities approved by Health Canada […]
    Output: equivalent
c. Input: The chief merchandising officer decides what the store is going to sell […]
    Output: equivalent
SST-5
a. Input: plodding, poorly written, murky and weakly acted, the picture feels as if everyone making it lost their movie mojo.
    Output: very negative
b. Input: duvall is strong as always .
    Output: very positive
c. Input: lohman adapts to the changes required of her , but the actress and director peter kosminsky never get the audience to break […]
    Output: neutral
MNLI
a. Input: This prosperous city has many museums, including a well-endowed Musee des Beaux-Arts (Square Verdrel) […]
    Output: False
b. Input: Duhame, who today makes her living as a graphic designer and illustrator, calls her book […]
    Output: Inconclusive
c. Input: At the agency or program level, it included management’s public commitment to reduce fraud and errors, as. Based on that information […]
    Output: True
DBpedia
a. Input: Lars Nielsen (born 3 November 1960 in Copenhagen) is a Danish rower.
    Output: athlete
b. Input: Calhoun County High School is a public secondary school in St. Matthews South Carolina USA.
    Output: educational institution
c. Input: David Goldschmid (sometimes credited as Dave Goldschmid) is an American television writer and producer currently writing for the daytime drama General Hospital.
    Output: artist
RTE
a. Input: In sub-Saharan Africa about one in every 30 people is infected with HIV.. 30% of the people infected with HIV live in Africa..
    Output: False
b. Input: The drawbacks of legalization do not imply that our current version of prohibition is the optimal drug strategy; it may well […]
    Output: False
c. Input: For example, the fields of Western farmers feed the United States and many other parts of the world, and India’s irrigation […]
    Output: True
HellaSwag
a. Input: The topic is Preparing salad. An illustrated egg, the website ”startcooking com” and ”vegetable salad” […]
    Output: is shown from above.
b. Input: The topic is Pets and Animals. [header] How to treat an injured rabbit’s paw [title] Identify sore hocks. [step] Pododermatitis […]
    Output: Once the condition has set in, though, you’ll need to take quick action to treat the injury. Leaving […]
c. Input: The topic is Playing squash. Two men stand on a racquetball court. the men
    Output: stretch then begin playing.
@@ C.2 Visualization of selected examples
Here we provide a umap (McInnes et al., 2018) visualization of selected examples. To avoid the denseness, we choose the annotation budget as 5. The visualization can be checked in Figure 6. First, comparing subfigures (a) and (b), we can clearly see that the selection of Vote-$k$ is much biased, and our IDEAL can identify a subset that is more favorable to be a proxy of full data. Second, comparing subfigures (c) and (d), we can see that the selected subset by Vote-$k$ is distributed on the right of full data. By comparison, our IDEAL can select a subset that is distributed more uniformly.
@@
Figure 6: Umap (McInnes et al., 2018) visualization to compare five selected examples from all examples using fully unsupervised methods Vote-$k$ and IDEAL (ours). Compared with Vote-$k$, IDEAL could choose the examples to better represent the whole data rather than get involved in diversity and including outliers.
@@ C.5 Prompt order in selective annotation
As pointed out by  (Lu et al., 2021), the performance of in-context learning is influenced not only by the selection of prompts but also by the order in which the prompts are presented to models. Although this work focuses solely on selective annotation problems, we are interested in exploring whether the selected subset can still lead to better performance when the order of prompts is permuted. Under an annotation budget of 18, we first retrieve prompts for each test instance from selected subsets achieved by different selective annotation methods. We then permute the order of prompts for each test instance 10 times, resulting in 10 different experimental trials. We show the average performance of these 10 trials and make a comparison between different selective annotation methods. We conduct experiments on MRPC, SST-5, and RTE datasets and present the results in Table 10. The results show that IDEAL outperforms baselines in 2 out of 3 cases, suggesting that our method can choose more stable and robust subsets against changed prompt orders.
@@ D.1 Details of datasets
In this paper, to demonstrate the superiority of our method, we employ 9 datasets typical datasets that have been widely used in previous NLP works (Su et al., 2023; Shi et al., 2023; Zhang et al., 2021) which can be categorized into 4 different tasks, including classification (MRPC (Dolan et al., 2004), SST-5 (Socher et al., 2013), MNLI (Williams et al., 2017), DBpedia (Lehmann et al., 2015), and RTE (Bentivogli et al., 2009)), multi-choice (HellaSwag (Zellers et al., 2019)), dialogue (MWoZ (Budzianowski et al., 2018)), and generation (GeoQuery (Zelle & Mooney, 1996) and Xsum (Narayan et al., 2018)). We list the datasets and the models used in Table 11.
Datasets Task Models Classification MRPC (Dolan et al., 2004) Paraphrase Detection GPT-Neo, GPT-J, GPT-3.5-Turbo SST-5 (Socher et al., 2013) Sentiment Analysis GPT-J DBpedia (Lehmann et al., 2015) Topic Classification GPT-J RTE (Bentivogli et al., 2009))   Natural Language Inference GPT-Neo, GPT-J, GPT-3.5-Turbo MNLI (Williams et al., 2017) Natural Language Inference GPT-Neo, GPT-J, GPT-3.5-Turbo Multiple-Choice HellaSwag (Zellers et al., 2019) Commonsense Reasoning GPT-J Dialogue MWoZ (Budzianowski et al., 2018) Dialogue State Tracking Text-davinci-002 Generation GeoQuery (Zelle & Mooney, 1996) Semantic Parsing Text-davinci-002 Xsum (Narayan et al., 2018) Summarization GPT-J

### v3

@@ 4.3.2 Comparisons with alternative methods
We also compare our method with other alternative methods that can select the coreset from large-scale unlabeled data. We perform the evaluations on MRPC, MNLI and HellaSwag. We include the following alternative methods (1) $K$-Means [27], which groups all examples into $m$ clusters, and selects the centroid example from each cluster. (2) Maximizing facility location (MFL) [28], which aims at optimizing the representativeness of the selected subset. (3) Fast Vote-$k$ [7], which is an efficient alternative to Vote-$k$ which directly picks $m$ examples with the largest Vote-$k$ scores.
@@
Method Datasets
Selection Retrieval MRPC MNLI HellaSwag
Vote-$k$ Similar 64.6 38.9 67.9
IDEAL Similar 66.4 41.0 68.6
Vote-$k$ Random 60.7 37.8 64.6
IDEAL Random 62.5 39.0 66.8
@@
Method Models Test Domain
IMDb BoolQ Cst.
Vote-$k$ GPT-Neo 71.1 56.4
IDEAL GPT-Neo 72.2 58.0
Vote-$k$ GPT-J 76.4 56.1
IDEAL GPT-J 76.8 56.4
@@ 4.3.3 Evaluation with different retrieval methods
In previous experiments, we used a similarity-based prompt retrieval method by default. In this section, we conduct experiments to quantify the effect of different prompt retrieval methods under the annotation 100. We present the results in Table 4. We observe that both Vote-$k$ and IDEAL suffer from a significant performance drop when the prompt retrieval method is changed from similarity-based to random selection. Notably, IDEAL also achieves better performance than Vote-$k$ when combined with random retrieval in all datasets. It suggests that IDEAL can cultivate a more stable training subset [29] for in-context learning tasks. Note that we also show that our IDEAL is more stable and robust against the order of in-context examples in Appendix C.5.
@@ 4.3.4 Evaluation on other language models
Here we evaluate IDEAL on other language models, including GPT-Neo 2.7B [25], and the advanced chat model GPT-3.5-Turbo. While GPT-3.5-Turbo has mainly been optimized for chat, it also performs well on traditional completion tasks [30]. To conduct experiments, we select three classification tasks (MRPC, MNLI, and RTE), considering they are easier for prompting GPT-3.5-Turbo to return responses without pleasantries or explanatory content.
The evaluation results are presented in Figure 5. Our evaluations reveal that IDEAL consistently outperforms the baselines across all models tested. This demonstrates the versatility of our method in the context of learning tasks using models of varying sizes. Notably, we observe that the largest model, i.e., GPT-3.5-Turbo, performs worse than GPT-Neo and GPT-J. This situation arises because GPT-3.5-Turbo is primarily optimized for chat tasks and faces challenges in following human instructions for classification. This scenario also has been identified in [31].
@@ 4.3.5 Evaluation on out-of-distribution tasks
We further evaluate our method on out-of-distribution tasks [32, 33, 34, 35, 36], where there is a distribution shift between the selective annotation data and test data. Following [29], we compare IDEAL and Vote-$k$ using SST-2 [10], BoolQ [37] as source tasks, and IMDb [38], BoolQ Contrast Set [39] as target tasks, respectively. In all evaluations, we set the annotation budget as 18 and use the similarity-based retrieve to perform the evaluations on the test set in target domains. We use GPT-J 6B and GPT-Neo 2.7B here and show the results in Table 4. We can observe that IDEAL still outperforms baselines on all datasets with two models, implying that IDEAL could select the subset which could depict the invariant properties of this kind of tasks and generalize to out-of-distribution scenarios.
@@ 4.4 Case study: automatic annotation
In previous experiments, we used a small set of manually annotated examples as candidate prompts to make predictions. Instead of this procedure, here we are interested in a case study that utilizes the subset selected by IDEAL to annotate all available unlabeled data automatically, leading to a larger set of candidate prompts. Specifically, we first choose an initial subset from the pool of unlabeled data using IDEAL and manually label this selected subset. Afterward, we simulate the information diffusion process from the initial subset to all other data, where we employ those activated data as prompts to predict upcoming activated data at each step and label them accordingly with prediction results. This process ultimately makes a fully labeled training dataset.
@@
Finally, all examples (including manual labeling and automatic labeling ) are utilized as potential prompts in conjunction with the prompt retrieve technique for final testing. We name this paradigm as Auto-IDEAL and compare it with Vote-$k$ and origin IDEAL on all classification datasets. We choose 300 training data for each dataset to perform experiments. The manual annotation budget is set to 150, i.e., half of the labels of the candidate prompts in Auto-IDEAL are annotated automatically.
Experimental results are provided in Table 5. As can be observed, Auto-IDEAL even achieves better performance than IDEAL in 4 of 5 cases. Notably, although the performance is worse on MNLI, it is still competitive (better than Vote-$k$). It suggests that expanding the candidate prompts through automatic annotation following the diffusion process can further boost the performance of IDEAL. It benefits from the fact that information only diffuses between similar examples. Therefore, unlabeled examples will be automatically annotated using the most similar annotated examples as prompts leading to a promising annotation success rate.
@@ A.1 In-context learning
In-context learning (ICL) has become a new paradigm for natural language processing (NLP), where large language models make predictions only based on contexts augmented with a few examples [40, 41, 42, 43, 44]. A series of works attempts to revise, enhance, and understand ICL, which include but are not limited to prompt tuning [45, 46, 47], analyzing intrinsic mechanism [48, 49, 50, 51], evaluations [52, 53], applications in multiple domains [54, 55, 56], and etc. Different from them, this paper studies selective annotations for ICL, which can effectively reduce the annotation cost in ICL. Furthermore, compared with recent work [7], as discussed in the main paper, this work is superior in many aspects, such as the end-to-end manner, mitigation of the trade-off between diversity and representativeness, theoretical guarantees, and better empirical performance.
@@ A.2 Coreset selection
Coreset selection focuses on selecting a small but highly informative subset from a large dataset for follow-up tasks, which can significantly reduce the data storage cost and training consumption [57, 58, 59, 60]. Most of the works on coreset selection target the scenes of supervised learning and classification [61, 62, 63]. Only a few works extend coreset selection into unsupervised cases [60, 7]. This paper studies unsupervised data selection for annotations in ICL, which reduces the annotation expenses of prompts and helps large language models become better few-shot learners. Also, it enjoys theoretical support. Therefore, this work is different from previous efforts and contributes to the research community.
@@ A.3 Data distillation
Data distillation [64, 65, 66, 67, 68, 69] is an alternative approach for dataset compression and curation, which is inspired by knowledge distillation. Different from coreset selection, this series of works target synthesizing a small but informative dataset as an alternative to the original dataset. However, data distillation is criticized for only synthesizing a small number of data points due to computational source limitations [70, 71]. The performances of data distillation and data selection are therefore not compared directly. Besides, it is under-explored about how to perform data distillation in an unsupervised manner on natural language processing tasks. Based on this analysis, the data distillation strategy is not involved in empirical evaluations.
@@ B.1.1 Lemma 1
The proof is inspired by [72]. We first let
@@ C.1 Selected examples
Dataset Input
MRPC a. Input: The two Democrats on the five-member FCC held a press conference to sway opinion against […]     Output: not equivalent a. Input: The report shows that drugs sold in Canadian pharmacies are manufactured in facilities approved by Health Canada […]     Output: equivalent c. Input: The chief merchandising officer decides what the store is going to sell […]     Output: equivalent
SST-5 a. Input: plodding, poorly written, murky and weakly acted, the picture feels as if everyone making it lost their movie mojo.     Output: very negative b. Input: duvall is strong as always .     Output: very positive c. Input: lohman adapts to the changes required of her , but the actress and director peter kosminsky never get the audience to break […]     Output: neutral
MNLI a. Input: This prosperous city has many museums, including a well-endowed Musee des Beaux-Arts (Square Verdrel) […]     Output: False b. Input: Duhame, who today makes her living as a graphic designer and illustrator, calls her book […]     Output: Inconclusive c. Input: At the agency or program level, it included management’s public commitment to reduce fraud and errors, as. Based on that information […]     Output: True
DBpedia a. Input: Lars Nielsen (born 3 November 1960 in Copenhagen) is a Danish rower.     Output: athlete b. Input: Calhoun County High School is a public secondary school in St. Matthews South Carolina USA.     Output: educational institution c. Input: David Goldschmid (sometimes credited as Dave Goldschmid) is an American television writer and producer currently writing for the daytime drama General Hospital.     Output: artist
RTE a. Input: In sub-Saharan Africa about one in every 30 people is infected with HIV.. 30% of the people infected with HIV live in Africa..     Output: False b. Input: The drawbacks of legalization do not imply that our current version of prohibition is the optimal drug strategy; it may well […]     Output: False c. Input: For example, the fields of Western farmers feed the United States and many other parts of the world, and India’s irrigation […]     Output: True
HellaSwag a. Input: The topic is Preparing salad. An illustrated egg, the website ”startcooking com” and ”vegetable salad” […]     Output: is shown from above. b. Input: The topic is Pets and Animals. [header] How to treat an injured rabbit’s paw [title] Identify sore hocks. [step] Pododermatitis […]     Output: Once the condition has set in, though, you’ll need to take quick action to treat the injury. Leaving […] c. Input: The topic is Playing squash. Two men stand on a racquetball court. the men     Output: stretch then begin playing.
@@ C.2 Visualization of selected examples
Here we provide a umap [73] visualization of selected examples. To avoid the denseness, we choose the annotation budget as 5. The visualization can be checked in Figure 6. First, comparing subfigures (a) and (b), we can clearly see that the selection of Vote-$k$ is much biased, and our IDEAL can identify a subset that is more favorable to be a proxy of full data. Second, comparing subfigures (c) and (d), we can see that the selected subset by Vote-$k$ is distributed on the right of full data. By comparison, our IDEAL can select a subset that is distributed more uniformly.
@@
Figure 6: Umap [73] visualization to compare five selected examples from all examples using fully unsupervised methods Vote-$k$ and IDEAL (ours). Compared with Vote-$k$, IDEAL could choose the examples to better represent the whole data rather than get involved in diversity and including outliers.
@@ C.5 Prompt order in selective annotation
As pointed out by  [74], the performance of in-context learning is influenced not only by the selection of prompts but also by the order in which the prompts are presented to models. Although this work focuses solely on selective annotation problems, we are interested in exploring whether the selected subset can still lead to better performance when the order of prompts is permuted. Under an annotation budget of 18, we first retrieve prompts for each test instance from selected subsets achieved by different selective annotation methods. We then permute the order of prompts for each test instance 10 times, resulting in 10 different experimental trials. We show the average performance of these 10 trials and make a comparison between different selective annotation methods. We conduct experiments on MRPC, SST-5, and RTE datasets and present the results in Table 10. The results show that IDEAL outperforms baselines in 2 out of 3 cases, suggesting that our method can choose more stable and robust subsets against changed prompt orders.
@@ D.1 Details of datasets
In this paper, to demonstrate the superiority of our method, we employ 9 datasets which can be categorized into 4 different tasks, including classification (MRPC [75], SST-5 [10], MNLI [76], DBpedia [17], and RTE [77]), multi-choice (HellaSwag [78]), dialogue (MWoZ [18]), and generation (GeoQuery [21] and Xsum [19]). We list the datasets and the models used in Table 11.
Datasets Task Models Classification MRPC [75] Paraphrase Detection GPT-Neo, GPT-J, GPT-3.5-Turbo SST-5 [10] Sentiment Analysis GPT-J DBpedia [17] Topic Classification GPT-J RTE [77])   Natural Language Inference GPT-Neo, GPT-J, GPT-3.5-Turbo MNLI [76] Natural Language Inference GPT-Neo, GPT-J, GPT-3.5-Turbo Multiple-Choice HellaSwag [78] Commonsense Reasoning GPT-J Dialogue MWoZ [18] Dialogue State Tracking Text-davinci-002 Generation GeoQuery [21] Semantic Parsing Text-davinci-002 Xsum [19] Summarization GPT-J